OpenStudy (anonymous):

Pls help me, i really don't understand this!??? Consider the equation 4x + 6y = 48. Part 1: On your own paper, graph this equation using the slope-intercept method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. Part 2: On your own paper, graph this equation using the intercepts method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar.

OpenStudy (anonymous):

y=mx + b (slope formula) m being the slope and b being the y - intercept. solve for y. 6y= 48-4x (subtract 4x to both sides) y= 8-2/3 x (divide 6 to both sides to get y) slope is -2/3 x y intercept is 8. plot 8 on the y intercept and from there count 2 downwards and 3 to the right. (rise/run)

OpenStudy (anonymous):

thank u sooooooooo much!!!